a car travels 30 meter for 30 minutes. the car then stop at a service station fir 10 minutes the car continues for another 20 minutes at a constant speed of 45m/s. draw the distance -time graph fir the car?
1
Expert's answer
2020-04-17T10:17:22-0400

A car travels 30 kilometers for 30 minutes. the car then stop at a service station fir 10 minutes the car continues for another 20 minutes at a constant speed of 45m/s. draw the distance -time graph fir the car?"45 \frac{m}{c} = 45\times3.6=162\frac{km}{h} ; S = v\times t=162\frac{km}{h}\times \frac{1}{3}=54;54+30=84"

all points are found